# MoRE: Multi-Modal Contrastive Pre-training with Transformers on X-Rays, ECGs, and Diagnostic Report

### v1

@@ Appendix B Pretraining Data Detail
The Mimic-IV ECG Matched Subset dataset Gow et al. (2023) also released in Physionet. This dataset is derived from the larger Mimic-IV Clinical dataset which is also the parent of the Mimic-CXR dataset. The dataset includes 800,000 ECG from 160,000 patients. The ECG dataset contains no labels but includes clinical text.
@@
Clinical Notes: For the clinical notes of Xray, we use text under the headings in format of ”Impression:”, and ”Findings:”. We filter out the text from these headings and remove any special characters and redundant spaces from the text. For ECG, there are multiple reports for a ECG data, we merge the first 7 reports as we find them to be available for most ECGs and contain the most information. We then remove any special characters and redundant spaces from the text.
@@ Appendix C Datasets
Pre-training Data For pretraining, we use matched subset from Mimic-CXR v2 Johnson et al. (2019) and Mimic-IV ECG Gow et al. (2023). Mimic-CXR v2 has about 224k frontal chest X-ray images and Radiology Report, and Mimic-IV ECG has about 800k ECG signals with short Cardiology Report. There are about 45k matched subset of X-ray and ECG from the same patients.
Datasets for Downstream Tasks
@@ Appendix D Data Pre-processing and Transformations
Pre-processing is vital to any deep learning model training as it can significantly impact the training times, convergence, and outcome. Especially for medical data, it is key to ensure no important features are lost and conversely, highlight the important features.
@@
ECG: We follow a few key pre-processing steps for ECG. 1.Re-Sampling: We resample the ECG data from 500Hz to 100Hz, changing its channel dimension from 5000 to 1000, making it a matrix of 12,1000. 2.Remove Nan: We find there are NA values present in the data so we swap any NA values with 0. 3.Baseline Wander: Baseline Wander is one of the pre-processing steps that absolutely cannot be missed. Xu et al. (2017) Baseline Wander is a low-frequency noise that is caused by the movement of the ECG leads. If not addressed can cause the model to deviate from the understanding of the data. Finally, we do 4.Per Lead Normalization with MinMax Scaling to bring the range of each lead to -1 and 1. For Transformation, we follow work of Raghu et al. (2022) and utilize two augmentations for variability, Time Warping: warp 4 segments of ECG by factor of 0.25, and Random Permutation: permute 4 segments of ECG in random order.
@@ Appendix E Implementation Detail
Our implementation utilizes the ‘ViT-Base-patch16-224’ model, pre-trained on ImageNet from the Timm library Wightman (2019), chosen for its ease of customization compared to the Hugging Face implementation. Our ViT base encoder consists of 12 transformer layers and 12 multi-head self-attention heads. We do not include any bias for query, key, value (qkv) during the linear projection of patches due to batch normalization after each layer. For the ECG modality, we employ a custom patch embedding strategy using two 1D-convolution layers, each followed by batch normalization and ReLU activation. The remainder of the ViT architecture is unchanged. The diagnostic reports are processed using ClinicalBert’s tokenizer with the parameter ‘add special tokens’ $=$ True to insert separator tokens between texts from the two modalities. We set ‘max length’ $=$ 512 and ‘truncation’ $=$ True. Our projection layer comprises two linear layers with a hidden dimension of 768 and an output dimension of 128. The first linear layer is bias-free, followed by batch normalization and ReLU activation, while the second layer includes a bias. We configure the temperature parameter of the InfoNCE loss van den Oord et al. (2018) at 0.1, making it learnable to optimize performance. The model is trained over 50 epochs with an early stopping criterion set at a patience of 10. We use the AdamW optimizer Zhuang et al. (2022), with a weight decay of 0.1 for pretraining and 0.02 for fine-tuning. We fine-tune only the projector layer and classifier head for linear evaluation and fine-tune the ’qkv’ weight of the last few transformer self attention layers for downstream tasks. Training leverages Automatic Mixed Precision (AMP) in PyTorch for enhanced speed and efficiency with minimal accuracy loss Micikevicius et al. (2017). Instead of using MoCo He et al. (2020) for larger batch sizes, we implement gradient accumulation over 4 steps, which has been shown to be effective in contrastive learning by simulating larger batch sizes Gao et al. (2023). The initial batch size is set at 100, and the model is trained on a single A100-SXM4-80GB GPU.

### v2

@@ Appendix B Pretraining Data Detail
Mimic-IV ECG Matched Subset dataset Gow et al. (2023) released in Physionet including 800,000 ECG from 160,000 patients. The dataset is derived from the larger Mimic-IV Clinical dataset which is also the parent of the Mimic-CXR dataset. The dataset contains no labels but includes clinical text.
@@
Clinical Notes:
For the clinical notes of Xray, we use text under the headings in format of ”Impression:”, and ”Findings:”. We filter out the text from these headings and remove any special characters and redundant spaces from the text. For ECG, there are multiple reports for a ECG data, we merge the first 7 reports as we find them to be available for most ECGs and contain the most information. We then remove any special characters and redundant spaces from the text.
@@ Appendix C Datasets
Pre-training Data For pretraining, we use matched subset from Mimic-CXR v2 Johnson et al. (2019) and Mimic-IV ECG Gow et al. (2023). Mimic-CXR v2 has about 224k frontal chest X-ray images and Radiology Report, and Mimic-IV ECG has about 800k ECG signals with short Cardiology Report. There are about 45k matched subset of X-ray and ECG from the same patients. Datasets for Downstream Tasks
@@ Appendix D Data Pre-processing and Transformations
Pre-processing is vital to deep learning model training as it can significantly impact the training times, convergence, and outcome. For medical data, it is key to ensure no important features are lost and conversely, highlight the important features.
@@
ECG:
We follow a few key pre-processing steps for ECG. 1.Re-Sampling: We resample the ECG data from 500Hz to 100Hz, changing its channel dimension from 5000 to 1000, making it a matrix of 12,1000. 2.Remove Nan: We find there are NA values present in the data so we swap any NA values with 0. 3.Baseline Wander: Baseline Wander is one of the pre-processing steps that absolutely cannot be missed. Xu et al. (2017) Baseline Wander is a low-frequency noise that is caused by the movement of the ECG leads. If not addressed can cause the model to deviate from the understanding of the data. Finally, we do 4.Per Lead Normalization with MinMax Scaling to bring the range of each lead to -1 and 1. For Transformation, we follow work of Raghu et al. (2022) and utilize two augmentations for variability, Time Warping: warp 4 segments of ECG by factor of 0.25, and Random Permutation: permute 4 segments of ECG in random order.
@@ Appendix E Implementation Detail
Our implementation utilizes the ‘ViT-Base-patch16-224’ model, pre-trained on ImageNet from the Timm library Wightman (2019), chosen for its ease of customization compared to the Hugging Face implementation. Our ViT base encoder consists of 12 transformer layers and 12 multi-head self-attention heads. We do not include any bias for query, key, value (qkv) during the linear projection of patches due to batch normalization after each layer. For the ECG modality, we employ a custom patch embedding strategy using two 1D-convolution layers, each followed by batch normalization and ReLU activation. The remainder of the ViT architecture is unchanged. The diagnostic reports are processed using ClinicalBert’s tokenizer with the parameter ‘add special tokens’ $=$ True to insert separator tokens between texts from the two modalities. We set ‘max length’ $=$ 512 and ‘truncation’ $=$ True. Our projection layer comprises two linear layers with a hidden dimension of 768 and an output dimension of 128. The first linear layer is bias-free, followed by batch normalization and ReLU activation, while the second layer includes a bias. We configure the temperature parameter of the InfoNCE loss van den Oord et al. (2018) at 0.1, making it learnable to optimize performance. The model is trained over 50 epochs with an early stopping criterion set at a patience of 10. We use the AdamW optimizer Zhuang et al. (2022), with a weight decay of 0.1 for pretraining and 0.02 for fine-tuning. We fine-tune only the projector layer and classifier head for linear evaluation and fine-tune the ’qkv’ weight of the last few transformer self attention layers for downstream tasks. Training leverages Automatic Mixed Precision (AMP) in PyTorch for enhanced speed and efficiency with minimal accuracy loss Micikevicius et al. (2017). Instead of using MoCo He et al. (2020) for larger batch sizes, we implement gradient accumulation over 4 steps, which has been shown to be effective in contrastive learning, simulating larger batch sizes Gao et al. (2023). The initial batch size is set at 100, the model is trained on a single A100-SXM4-80GB GPU.